OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation y - 4x = -3 slope = -1; y-intercept = (0, -3) slope = -1 fourth; y-intercept = (0, 3) slope = 4; y-intercept = (0, -3) slope = 4; y-intercept = (0, 3)

OpenStudy (anonymous):

y=4x-3 slope m=4 y intercept=(0,-3)

OpenStudy (anonymous):

To make it easier try putting it in slope intercept form: y=mx+b

OpenStudy (anonymous):

@sai1234 Try explaining that. Dont just give answers.

OpenStudy (anonymous):

ok y=4x-3 it is in slope intercept form: y=mx+b slope m=4 y intercept (c)=0,-3)